OpenStudy (anonymous):

The length of the hypotenuse of a 30°-60°-90° triangle is 13 meters. Find the length of the side opposite the 30° angle. a. 6.5 m c. 11.3 m b. 26.0 m d. 22.5 m

OpenStudy (anonymous):

@Thehulk49

OpenStudy (anonymous):

You have to use sin30: opposite length/hypotenuse Sin30: 0.5 So 0.5=opposite lenght/13 Opposite lenght=13x0.5 Opposite lenght: 6.5 Latter a

OpenStudy (anonymous):

omg thank yaaaaaaallllllllll c':

OpenStudy (anonymous):

This is what I got. The length of the hypotenuse of a 30-60-90 triangle is 13 meters. Find the length of the opposite side 30 degree angle. sohcahtoa: sin = opposite(opp)/hypotenuse(hyp), cos = adjacent(adj)/hyp, tan = opp/adj a 30-60-90 triangle is half an equilateral triangle if the equilateral triangle has sides 1,1, and 1 then: sin 30 = 0.5/1 = 0.5 (0.5 is half of 1) cos 30 = adj/1 --> adj = sqrt(3)/2 30-60-90 triangles have the ratio 0.5:sqrt(3)/2:1 as their sides so if the hypotenuse is 13 meters then the side opposite the 30 degree angle is: 13 * 0.5 = 6.5 meters

OpenStudy (anonymous):

So the answer is 6.5